OpenStudy (anonymous):

Express in simplest form 10 square root of 24 divide 2 square root of 2?

OpenStudy (mathlegend):

\[\frac{ 10\sqrt{24} }{ 2\sqrt{2} }\]

OpenStudy (anonymous):

yeah

OpenStudy (mathlegend):

So what two numbers can you multiply together, one being a perfect square that goes into 24? a x b = 24 a or b should be a perfect square

OpenStudy (anonymous):

4 and 6

OpenStudy (mathlegend):

Good

OpenStudy (mathlegend):

So, now what is the square root of 4?

OpenStudy (anonymous):

2

OpenStudy (mathlegend):

So that comes to the outside... 10(2)\[10(2)\sqrt{6}\]

OpenStudy (anonymous):

so that would be the answer

OpenStudy (mathlegend):

\[20\sqrt{6}\]

OpenStudy (mathlegend):

\[\frac{ 20\sqrt{6} }{ 2\sqrt{2} }\]

OpenStudy (anonymous):

Okay thanks now I get it

OpenStudy (mathlegend):

So what is your final answer @farmergirl411

OpenStudy (anonymous):

10 square root of 3

OpenStudy (mathlegend):

\[\frac{ 10\sqrt{6} }{ \sqrt{2}}\]

OpenStudy (mathlegend):

I forget if we can actually simplify those two radicals like that... let me double check

OpenStudy (anonymous):

ok

OpenStudy (mathlegend):

Oh okay I just checked and it is \[10\sqrt{3}\]

OpenStudy (anonymous):

ok thanks

hero (hero):

\[\frac{ 10\sqrt{24} }{ 2\sqrt{2}} \\=\frac{ 10\sqrt{4 \times 6} }{ 2\sqrt{2}} \\=\frac{ 10\sqrt{4} \times \sqrt{6}}{ 2\sqrt{2}} \\= \frac{ 10(2)\sqrt{6}}{ 2\sqrt{2}} \\= 10\frac{\sqrt{6}}{\sqrt{2}} \\= 10\sqrt{\frac{6}{2}} \\= 10\sqrt{3}\]

OpenStudy (mathlegend):

Yeah, see that looks better... it was confusing me because I had the 6 and 2 under separate radicals...

OpenStudy (mathlegend):

I think I was thinking of addition...

OpenStudy (mathlegend):

or subtraction... you know?

hero (hero):

Remember this rule \[\frac{\sqrt{a}}{\sqrt{b}} = \sqrt{\frac{a}{b}}\]

OpenStudy (mathlegend):

Will do! :D Thanks again @Hero